PLEASE PLEASE BE SURE THE ANSWER IS RIGHT, THANKSRequired information A 1.5-in. steel plate and a 0.75-in. gray cast iron plate are compressed with one bolt and nut. The bolt
Answer #1

Solution Gaiß length l75+0.75 +27(0,095) d=2.44 in for I Ň washer For Material stiffness we draw following foustrum ki R 10.K2=31-73 Mebflin ki = 0:5774 30 x0.531 In (1.15580.095 +0.75-0.531) (0.7570.53), (1155x0.098 +0.75 +0.531) (0.75-01531) Ki=89for Gray Cast iron foutum (kq) it = 0.75 d=0,5 D=0.75+2x0.095 to 30% E = 14.5 MPsi D=0.860 hence Kq=0.5774 T (145)*0.5 In 1-1